THE STUDY OF INTERMOLECULAR INTERACTIONS IN LIQUID CHLOROFORM AND ITS SOLU-TIONS BY RAMAN SPECTRA

F. H. TUKHVATULLIN, \underline{A. JUMABAEV}, U. N. TASHKENBAEV, Samarkand State University, 703004 Samarkand, Univ. Blvd, 15, Uzbekistan; B. S. OSMANOV, Department of Physics, Ohio State University, Columbus OH 43210.